# Understanding the Cluster LP for Correlation Clustering

### v2

@@ Abstract
In the classic Correlation Clustering problem introduced by Bansal, Blum, and Chawla [BBC04], the input is a complete graph where edges are labeled either $+$ or --, and the goal is to find a partition of the vertices that minimizes the sum of the $+$edges across parts plus the sum of the --edges within parts. In recent years, Chawla, Makarychev, Schramm and Yaroslavtsev [CMSY15] gave a 2.06-approximation by providing a near-optimal rounding of the standard LP, and Cohen-Addad, Lee, Li, and Newman [CLN22, CLLN23] finally bypassed the integrality gap of 2 for this LP giving a $1.73$-approximation for the problem.
@@
In order to create a simple and unified framework for Correlation Clustering similar to those for typical approximate optimization tasks, we propose the cluster LP as a strong linear program that might tightly capture the approximability of Correlation Clustering. It unifies all the previous relaxations for the problem. It is exponential-sized, but we show that it can be $(1+\varepsilon)$-approximately solved in polynomial time for any $\varepsilon>0$, providing the framework for designing rounding algorithms without worrying about correlated rounding errors; these errors are handled uniformly in solving the relaxation.
We demonstrate the power of the cluster LP by presenting a simple rounding algorithm, and providing two analyses, one analytically proving a $1.49$-approximation and the other solving a factor-revealing SDP to show a $1.437$-approximation. Both proofs introduce principled methods by which to analyze the performance of the algorithm, resulting in a significantly improved approximation guarantee.
Finally, we prove an integrality gap of $4/3$ for the cluster LP, showing our $1.437$-upper bound cannot be drastically improved. Our gap instance directly inspires an improved NP-hardness of approximation with a ratio $24/23\approx 1.042$; no explicit hardness ratio was known before.
@@ 1 Introduction
Clustering is a classic problem in unsupervised machine learning and data mining. Given a set of data elements and pairwise similarity information between the elements, the task is to find a partition of the data elements into clusters to achieve (often contradictory) goals of placing similar elements in the same cluster and separating different elements in different clusters. Introduced by Bansal, Blum, and Chawla [BBC04], Correlation Clustering elegantly models such tension and has become one of the most widely studied formulations for graph clustering. The input of the problem consists of a complete graph $(V,E^{+}\uplus E^{-})$, where $E^{+}\uplus E^{-}={V\choose 2}$, $E^{+}$ representing the so-called positive edges and $E^{-}$ the so-called negative edges. The goal is to find a clustering (partition) of $V$, namely $(V_{1},\dots,V_{k})$, that minimizes the number of unsatisfied edges, namely the $+$edges between different clusters and the --edges within the same cluster. Thanks to the simplicity and modularity of the formulation, Correlation Clustering has found a number of applications, e.g., finding clustering ensembles [BGU13], duplicate detection [ARS09], community mining [CSX12], disambiguation tasks [KCMNT08], automated labelling [AHK+09, CKP08] and many more.
@@
In this work, we propose the cluster LP as a single relaxation that captures all of the existing algorithmic results. Based on this new unified framework, we design a new rounding algorithm as well as principled tools for the analysis that significantly extend the previous ones, ultimately yielding a new approximation ratio of $1.437+\varepsilon$. The study of the cluster LP sheds light on the hardness side as well, as we prove a $4/3\approx 1.33$ gap for the cluster LP and a $24/23\approx 1.042$ NP-hardness of approximation.
@@ 1.1 Our Results
The cluster LP is the most powerful LP that has been considered for the problem. Indeed, previous algorithms in [CLN22] and [CLLN23] can be significantly simplified if one is given a $(1+\varepsilon)$-approximate solution to the LP. A large portion of the algorithms and analysis in [CLN22] and [CLLN23] is devoted to handle the additive errors incurred by the correlated rounding procedure, which is inherited from the Raghavendra-Tan rounding technique [RT12]. Instead, we move the complication of handling rounding errors into the procedure of solving the cluster LP relaxation.
With this single powerful relaxation, we believe that Theorem 1 provides a useful framework for future work that may use more ingenious rounding of the exponential-sized cluster LP without worrying about errors. Indeed, the constraints in the cluster LP imply that the matrix $(1-x_{uv})_{u,v\in V}$ is PSD,  [^6] Consider the matrix $Y\in[0,1]^{V\times V}$ where $y_{uv}=1-x_{uv}$ for every $u,v\in V$ ($Y_{uu}=1,\forall u\in V$). For every $w\in\mathbb{R}^{V}$, we have $w^{T}Yw=\sum_{u,v\in V}y_{uv}w_{u}w_{v}=\sum_{u,v}\sum_{S\supseteq\{u,v\}}z_{S}w_{u}w_{v}=\sum_{u,v}\sum_{S\subseteq V}z_{S}\cdot(w_{u}\cdot 1_{u\in S})\cdot(w_{v}\cdot 1_{v\in S})=\sum_{S\subseteq V}z_{S}\left(\sum_{u\in S}w_{u}\right)\left(\sum_{v\in S}w_{v}\right)\geq 0$. and thus the LP is at least as strong as the natural SDP for the problem. For the complementary version of maximizing the number of correct edges, the standard SDP is known to give a better approximation guarantee of $0.766$ [Swa04, CGW05]. For the minimization version, the standard SDP has integrality gap at least 1.5 (see Appendix C), but it is still open whether this program has an integrality gap strictly below 2 or not.
@@
There exists a $(1.49+\varepsilon)$-approximation algorithm for Correlation Clustering that runs in time $O(n^{{\mathrm{poly}}(1/\varepsilon)})$.
@@
In order to obtain an even tighter analysis of the same algorithm, we introduce the new factor revealing SDP that searches over possible global distributions of triangles in valid Correlation Clustering instances. By numerically solving such an SDP, we can further improve the approximation ratio of the same algorithm.
@@
There exists a $(1.437+\varepsilon)$-approximation algorithm for Correlation Clustering that runs in time $O(n^{{\mathrm{poly}}(1/\varepsilon)})$.
@@ 1.2 Further Related Work
2 Algorithmic Framework and Setup for Analysis
@@
In this section, we describe our algorithm for obtaining the improved approximation ratio for Correlation Clustering. We solve the cluster LP using Theorem 1 to get a fractional solution $z=(z_{S})_{S\subseteq V}$, which determines $x\in[0,1]^{V\choose 2}$ as in (2): $x_{uv}:=1-\sum_{S\supseteq\{u,v\}}z_{S}$ for every $uv\in{V\choose 2}$. We have $\mathrm{obj}(x)\leq(1+\varepsilon){\mathrm{opt}}$. The theorem will be proved in Section 4. With $z$, we then run two procedures: the cluster-based rounding and the pivot-based rounding with threshold $1/3$. We select the better result as the final clustering. The two procedures are defined in Algorithms 1 and 2 respectively. We use $N^{+}(u)$ and $N^{-}(u)$ to denote the sets of $+$ and --neighbors of a vertex $u\in V$ respectively.
@@
Algorithm 2 Pivot-Based Rounding with Threshold $1/3$
1:${\mathcal{C}}\leftarrow\emptyset,V^{\prime}\leftarrow V$
@@
3:     randomly choose a pivot $u\in V^{\prime}$
4:     $C\leftarrow\{v\in V^{\prime}\cap N^{+}(u):x_{uv}\leq\frac{1}{3}\}$
5:     for every $v\in V^{\prime}\cap N^{-}(u)$ do independently add $v$ to $C$ with probability $1-x_{uv}$
6:     randomly choose a set $S\ni u$, with probabilities $z_{S}$triangleright We have $\sum_{S\ni u}z_{S}=1$
7:     $C\leftarrow C\cup(S\cap V^{\prime}\cap N^{+}(u))$, ${\mathcal{C}}\leftarrow{\mathcal{C}}\cup\{C\}$, $V^{\prime}\leftarrow V^{\prime}\setminus C$
8:return ${\mathcal{C}}$
@@
Therefore, a $+$edge $uv$ will incur a cost of $\frac{2x_{uv}}{1+x_{uv}}$ in expectation in the cluster-based rounding procedure, and a --edge will incur a cost of $\frac{1-x_{uv}}{1+x_{uv}}$. The approximation ratios for a $+$edge $uv$ and a --edge $uv$ are respectively $\frac{2}{1+x_{uv}}$ and $\frac{1}{1+x_{uv}}$. Notice that the latter quantity is at most $1$.
@@
We now proceed to the pivot-based rounding procedure in Algorithm 2. We remark that to recover the correlated rounding algorithm in [CLN22] and [CLLN23], we can use $C\leftarrow\emptyset$ in Step 4. Then we can obtain their approximation ratios without the complication of handling rounding errors. The errors are handled in [CLN22] by distinguishing between the short, median and long $+$edges. In our algorithm, we also distinguish between short $+$edges (those with $x_{uv}\leq\frac{1}{3}$) and long $+$edges (those with $x_{uv}>\frac{1}{3}$); however, the purpose of this distinction is to get an improved approximation ratio, instead of to bound the rounding errors.
Our high-level setup of the analysis follows from [CLN22, CLLN23], which in turn is based on [ACN08] and [CMSY15]. We consider a general budget for every edge. We shall define two budget functions:
  • $b^{+}:[0,1]\to\mathbb{R}_{\geq 0}$ and $b^{-}:[0,1]\to\mathbb{R}_{\geq 0}$.
They determine the budget $b_{uv}$ for the edge $uv$: if $uv\in E^{+}$, then $b_{uv}:=b^{+}(x_{uv})$, and if $uv\in E^{-}$, then $b_{uv}:=b^{-}(x_{uv})$.
We now focus on one iteration of the while loop in Algorithm 2. Suppose $u,v,w\in V^{\prime}$ at the beginning of the iteration, and let $C$ be the cluster constructed at the end. We use $u$ to denote the event that $u$ is chosen as the pivot. We say $vw$ incurs a cost in the iteration, if $vw\in E^{+}$ and $|C\cap\{v,w\}|=1$, or $vw\in E^{-}$ and $\{v,w\}\subseteq C$. Then, we define
$\mathrm{cost}_{u}(v,w):=\Pr[vw\text{ incurs a cost}\mid u].$
@@
$\frac{1}{|V^{\prime}|}\sum_{u\in V^{\prime}}\sum_{(v,w)\in{V^{\prime}\choose 2}}\mathrm{cost}_{u}(v,w)\frac{1}{|V^{\prime}|}\sum_{(u,v,w)\in{V^{\prime}\choose 3}}\mathrm{cost}(u,v,w)+\frac{1}{|V^{\prime}|}\sum_{(u,v)\in{V^{\prime}\choose 2}}\mathrm{cost}(u,v).$
@@
To obtain an approximation ratio of $\alpha\in[1,2)$, we consider a variant of our algorithm, in which we run the cluster-based rounding procedure (Algorithm 1) with probability $\frac{\alpha}{2}$, and the pivot-based rounding procedure with threshold $1/3$ (Algorithm 2) with the remaining probability $1-\frac{\alpha}{2}$. Clearly, the actual algorithm that picks the better of the two clusterings generated can only be better. We set up the budget functions $b^{+}$ and $b^{-}$ such that every edge pays a cost of at most $\alpha$ times its LP cost in expectation. That is, the following properties are satisfied for every $x\in[0,1]$:
$\frac{\alpha}{2}\cdot\frac{2x}{1+x}+\left(1-\frac{\alpha}{2}\right)b^{+}(x)=\alpha x,\qquad\frac{\alpha}{2}\cdot\frac{1-x}{1+x}+\left(1-\frac{\alpha}{2}\right)b^{-}(x)=\alpha(1-x).$
@@
$b^{+}_{\alpha}(x):=\frac{\alpha}{1-\alpha/2}\cdot\frac{x^{2}}{1+x},\qquad\text{and}\qquad b^{-}_{\alpha}(x):=\frac{\alpha}{1-\alpha/2}\cdot\frac{(1+2x)(1-x)}{2(1+x)},\qquad\forall x\in[0,1].$ (5)
@@
If the budget functions $b^{+}_{\alpha}$ and $b^{-}_{\alpha}$ satisfy (4) for some $\alpha\in[1,2)$, then our algorithm has an approximation ratio of $\alpha$.

### v3

@@ Abstract
In the classic Correlation Clustering problem introduced by Bansal, Blum, and Chawla [BBC04], the input is a complete graph where edges are labeled either ++ or -, and the goal is to find a partition of the vertices that minimizes the sum of the ++edges across parts plus the sum of the -edges within parts. In recent years, Chawla, Makarychev, Schramm and Yaroslavtsev [CMSY15] gave a 2.06-approximation by providing a near-optimal rounding of the standard LP, and Cohen-Addad, Lee, Li, and Newman [CLN22, CLLN23] finally bypassed the integrality gap of 2 for this LP giving a $1.73$-approximation for the problem.
@@
In order to create a simple and unified framework for Correlation Clustering similar to those for typical approximate optimization tasks, we propose the cluster LP as a strong linear program for Correlation Clustering. It is exponential-sized, but we show that it can be $(1+\varepsilon)$-approximately solved in polynomial time for any $\varepsilon>0$, providing the framework for designing rounding algorithms without worrying about correlated rounding errors; these errors are handled uniformly in solving the relaxation.
We demonstrate the power of the cluster LP by presenting new rounding algorithms, and providing two analyses, one analytically proving a $1.56$-approximation and the other solving a factor-revealing SDP to show a $1.485$-approximation. Both proofs introduce principled methods by which to analyze the performance of the algorithm, resulting in a significantly improved approximation guarantee.
Finally, we prove an integrality gap of $4/3$ for the cluster LP, showing our $1.485$-upper bound cannot be drastically improved. Our gap instance directly inspires an improved NP-hardness of approximation with a ratio $24/23\approx 1.042$; no explicit hardness ratio was known before.
@@ 1 Introduction
Clustering is a classic problem in unsupervised machine learning and data mining. Given a set of data elements and pairwise similarity information between the elements, the task is to find a partition of the data elements into clusters to achieve (often contradictory) goals of placing similar elements in the same cluster and separating different elements in different clusters. Introduced by Bansal, Blum, and Chawla [BBC04], Correlation Clustering elegantly models such tension and has become one of the most widely studied formulations for graph clustering. The input of the problem consists of a complete graph $(V,E^{+}\uplus E^{-})$, where $^{+}\uplus E^{-}={V\choose 2}, E+E^{+}$ representing the so-called positive edges and $E^{-}$ the so-called negative edges. The goal is to find a clustering (partition) of $V$, namely $(V_{1},\dots,V_{k})$, that minimizes the number of unsatisfied edges, namely the ++edges between different clusters and the -edges within the same cluster. Thanks to the simplicity and modularity of the formulation, Correlation Clustering has found a number of applications, e.g., finding clustering ensembles [BGU13], duplicate detection [ARS09], community mining [CSX12], disambiguation tasks [KCMNT08], automated labelling [AHK+09, CKP08] and many more.
@@
In this work, we propose the cluster LP as a single relaxation that captures all of the existing algorithmic results. Based on this new unified framework, we design a new rounding algorithm as well as principled tools for the analysis that significantly extend the previous ones, ultimately yielding a new approximation ratio of $1.485+\varepsilon$. The study of the cluster LP sheds light on the hardness side as well, as we prove a $4/3\approx 1.33$ gap for the cluster LP and a $24/23\approx 1.042$ NP-hardness of approximation.
@@ 1.1 Our Results
The cluster LP is one of the most powerful LPs that has been considered for the problem. A large portion of the algorithms and analysis in [CLN22] and [CLLN23] is devoted to handle the additive errors incurred by the correlated rounding procedure, which is inherited from the Raghavendra-Tan rounding technique [RT12]. Instead, we move the complication of handling rounding errors into the procedure of solving the cluster LP relaxation.
With this single powerful relaxation, we believe that Theorem 1 provides a useful framework for future work that may use more ingenious rounding of the exponential-sized cluster LP without worrying about errors. Indeed, the constraints in the cluster LP imply that the matrix $(1-x_{uv})_{u,v\in V}$ is PSD,  [^7] Consider the matrix $Y\in[0,1]^{V\times V}$ where $y_{uv}=1-x_{uv}$ for every $u,v\in V$ ($Y_{uu}=1,\forall u\in V$). For every $w\in\mathbb{R}^{V}$, we have $w^{T}Yw=\sum_{u,v\in V}y_{uv}w_{u}w_{v}=\sum_{u,v}\sum_{S\supseteq\{u,v\}}z_{S}w_{u}w_{v}=\sum_{u,v}\sum_{S\subseteq V}z_{S}\cdot(w_{u}\cdot 1_{u\in S})\cdot(w_{v}\cdot 1_{v\in S})=\sum_{S\subseteq V}z_{S}\left(\sum_{u\in S}w_{u}\right)\left(\sum_{v\in S}w_{v}\right)\geq 0$. and thus the LP is at least as strong as the natural SDP for the problem. For the complementary version of maximizing the number of correct edges, the standard SDP is known to give a better approximation guarantee of $0.766$ [Swa04, CGW05]. For the minimization version, the standard SDP has integrality gap at least 1.5 (see Appendix A), but it is still open whether this program has an integrality gap strictly below 2 or not.
@@
There exists a $(1.56+\varepsilon)$-approximation algorithm for Correlation Clustering that runs in time $O(n^{{\mathrm{poly}}(1/\varepsilon)})$.
@@
In order to obtain an even tighter analysis of a similar algorithm, we introduce the new factor revealing SDP that searches over possible global distributions of triangles in valid Correlation Clustering instances. By numerically solving such an SDP, we can further improve the approximation ratio of the same algorithm.
@@
There exists a $(1.485+\varepsilon)$-approximation algorithm for Correlation Clustering that runs in time $O(n^{{\mathrm{poly}}(1/\varepsilon)})$.
@@ 1.2 Further Related Work
2 Overview and Algorithm
In this section, we provide an overview of our algorithms and tools for the analysis. Section 2.1 introduces some techniques used for formulating and solving the cluster LP. Section 2.2 presents our full rounding algorithm and provides high-level ideas for the analysis. Section 2.3 discusses ideas behind our gap and hardness results. These techniques were introduced over the course of three conference papers [CLN22, CLLN23, CCL+24], so we try to highlight their connections and developments over time. For instance, since the cluster LP was formalized in the last conference paper [CCL+24], several of the following techniques were initially presented in relation to a direct rounding of our initial (Sherali-Adams-based) relaxation.
2.1 Formulating and Solving Cluster LP
Previous Algorithms.
To understand in more detail our contribution and the new techniques introduced, we need to provide a brief summary of previous approaches. We first recall the classic LP relaxation (whose integrality gap is known to be in $[2,2.06]$). There is a variable $x_{uv}$ for each pair of vertices whose intended value is 1 if $u,v$ are not in the same cluster and 0 otherwise. The goal is thus to minimize $\sum_{uv\in E^{+}}x_{uv}+\sum_{uv\in E^{-}}(1-x_{uv})$, under the classic triangle inequality constraint: $\forall u,v,w$, $x_{uv}\leq x_{uw}+x_{wv}$.
To obtain the 2.5-approximation algorithm, Ailon, Charikar and Newman [ACN08] designed a pivot-based rounding method that proceeds as follows: (1) Pick a random vertex $p$, called the pivot, and create the cluster containing $p$; and (2) Recurse on the rest of the instance (the vertices not in the cluster). This hence builds the clustering in a sequential manner. Once we commit to this scheme, the main design question that remains is how to construct the cluster containing the pivot. The solution of [ACN08] was to go over all the other vertices and for each other vertex $u$ place it in the cluster of the pivot $p$ with probability $(1-x_{pu})$, independently of the other random decisions made for the other vertices.
This was further improved by Chawla, Makarychev, Schramm and Yaroslavtsev [CMSY15] who kept the same scheme but provided a better rounding approach: For ++edges, they replaced the probability $(1-x_{pu})$ with a new rounding function $f^{+}$ to apply to the quantity $(1-x_{pu})$ yielding the probability of incorporating $u$ in $p$’s cluster. Similar to the approach of [ACN08], the (streamlined) analysis of this rounding scheme was triangle-based: The analysis is a charging scheme that charges the cost paid by each pair $u,v$ to the triangle $p,u,v$ where $p$ is the pivot which decided edge $u,v$. Then the crux of the analysis is to show that for any triangle $p,u,v$, the charge is bounded compared to the LP cost.
Sherali-Adams and Correlated Rounding.
One potential place for an improvement in the previous algorithms is the independent rounding given a pivot $p$; in both [ACN08, CMSY15], once we choose $p$, other vertices are put into the cluster of $p$ independently of one another. It might be more advantageous to make these decisions in a correlated way. However, the standard LP, which only has variables $\{x_{uv}\}_{uv\in\binom{V}{2}}$ for the pairs, is not suited to provide information about such correlations.
Motivated by this fact, our first new idea, which appeared in the first conference version [CLN22], is the usage of correlated rounding based on the Sherali-Adams hierarchy. Namely, given a pivot $p$, the set of ++neighbors of $p$ that join the cluster of $p$ is chosen in a correlated manner, using the techniques of Raghavendra and Tan [RT12]. Concretely, the Sherali-Adams hierarchy provides variables of the form $y_{S}$ for any constant-sized set of vertices $S$ that indicates the probability that all the vertices in $S$ are in the same cluster. [^8] In which case, $y_{uv}=1-x_{uv}$. Given a pivot $p$, the correlated rounding then allows to sample in such a way that the probability that $u,v$ join the cluster of $p$ is $y_{puv}\pm\varepsilon$, where $\varepsilon$ is an arbitrarily small constant. Applying this rounding to a set of vertices $S$, it ensures that we can sample a subset $C$ of $S$ both being of arbitrary size (importantly not necessarily constant) such that pairs $u,v$ ends up in $C$ with probability $y_{puv}$ on average, up to losing an additive $\varepsilon|S|^{2}$ error in the cost.
The main issue we need to face with the above plan of attack is that the additive error of $\varepsilon|S|^{2}$, which we pay when we create a cluster $C$, may end up being much larger than the cost of the optimum solution. In fact, an optimum solution may have cost zero. The correlated rounding was initially developed for maximization problems like Max-CSPs [dlVM07, BRS11, GS11, RT12, YZ14] where this additive error can be easily ignored by only additively sacrificing $\varepsilon$ in the approximation ratio, but this is not the case for Correlation Clustering.  [^9] The only exception is [GS11] that studied minimization versions including Min Bisection and Min Uncut, but their technique heavily relies on the property of cuts not easily adaptable to Correlation Clustering.
[CLN22] applied this rounding to +neighbors of $p$ to get a $(2+\varepsilon)$-approximation (with respect to a $O(1/\varepsilon^{2})$-level of Sherali-Adams) and further combined it with the global triangle charging scheme to achieve a $(1.994+\varepsilon)$-approximation. This charging scheme was later further developed and used to analyze global triangle distributions, which is discussed at the end of Section 2.2.
Dealing with additive errors from the correlated rounding is one of the most technical parts [CLN22] and perhaps a limitation to getting an improved approximation ratio (e.g., special attention must be paid to edges with values close to 0 and $1$ in both the algorithm and in the analysis). That brings us to our second main idea of preclustering, so that we can (1) precluster the instance so that the cost of the optimal solution becomes relatively large compared to the size of the “unsure part”, and (2) use correlated rounding only for the unsure part of the instance so that the error from the correlated rounding is negligible.
Preclustering.
Preclustering is a subroutine that aims at identifying “clear clusters”. Formally, our preclustering outputs a pair $({\mathcal{K}},E_{\mathrm{adm}})$, where $\mathcal{K}$} is a partition of $V$ into so-called atoms and $E_{\mathrm{adm}}\subseteq{V\choose 2}$ is a set of admissible edges. The guarantee is that for any $\varepsilon>0$, there is a $(1+\varepsilon)$-approximate solution that is consistent with the preclustering; an atom is not broken, and if $u$ and $v$ are not in the same atom and $uv\notin E_{\mathrm{adm}}$, then $u$ and $v$ are in different clusters. We also ensure that $|E_{\mathrm{adm}}|\leq O_{\varepsilon}({\mathrm{opt}})$, so the number of edges on which we apply the correlated rounding (and pay the additive error) is comparable to $\mathrm{opt}$}.
The algorithm first constructs atoms. Intuitively, consider a cluster of the optimum (integral) solution that has $s$ vertices but contributes significantly smaller than $s^{2}$ to the objective function, say $\varepsilon s^{2}$. (Otherwise, an $\varepsilon^{2}s^{2}$-additive error will already be a $\varepsilon$-multiplicative error.) For this to happen, it must be that the number of -edges internal to $S$ is small compared to $s\choose 2$} and that the number ++edges with exactly one endpoint in $S$ is also much smaller than $s\choose 2$}. This cluster is thus almost a clique with tiny ++edge-expansion, which is easy to (approximately) identify. This idea was used by Cohen-Addad, Lattanzi, Mitrovic, Norouzi-Fard, Parotsidis, and Tarnawski for a massively-parallel algorithm with a large constant factor approximation [CLM+21]. Our first version of preclustering [CLLN23] used [CLM+21] as a black-box to construct atoms, but in this paper, we present the simplified version of [CCL+24], which starts from an arbitrary $O(1)$-approximate clustering and removes a vertex $v$ from a cluster $C$ when the ++neighborhood of $v$ is somewhat different from $C$ (and makes $v$ a singleton cluster).
The set of admissible edges is roughly defined as follows: we construct a graph $(V,E^{1})$ where two vertices are neighbors if their ++degrees are similar (up to a factor of $O(1/\varepsilon)$). Then an edge $uv$ is admissible if $u$ and $v$ have many (at least an $\Omega(\varepsilon)$ fraction) common neighbors in $E^{+}\cap E^{1}$. Here, the somewhat relaxed notions of degree-similarity and large-common-neighborhood (compared to the ones for the atoms, where we wanted ++neighborhood of $u$ and $v$ to be almost identical for $u$ and $v$ to be in the same atom) ensure that there is a near-optimal solution respecting the admissible edges, while also guaranting that the number of the admissible edges (crossing atoms) is not too large compared to $\mathrm{opt}$}.
Solving Cluster LP on Preclustered Instance by Sampling.
After constructing a preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$, we formulate an LP relaxation aimed at finding a $(1+\varepsilon)$-approximate good clustering for $({\mathcal{K}},E_{\mathrm{adm}})$, that we call the bounded sub-cluster LP. First introduced in [CLLN23], it is a strengthened version of the Sherali-Adams hierarchy used in [CLN22]. With a solution $(x,y)$ to the LP, we run a procedure that constructs a single cluster $C$ randomly. The probability that any vertex is in $C$ is precisely $1/y_{\emptyset}$, where $y_{\emptyset}$ is the fractional number of clusters in $y$. The probabilities that exactly one of $u$ and $v$ is in $C$, and both of them are in $C$, are respectively $\frac{x_{uv}}{y_{\emptyset}}$ and $\frac{1-x_{uv}}{y_{\emptyset}}$ up to some error terms arising from the correlated rounding. As usual, $x_{uv}$ is the extent in which $u$ and $v$ are separated.
To construct the solution $z=(z_{S})_{S\subseteq V}$ for the cluster LP, we generate $y_{\emptyset}\Delta$ many clusters $C$ independently, for a large enough polynomial $\Delta$. Roughly speaking, the solution $z$ is $\frac{1}{\Delta}$ times the multi-set of clusters $C$ we generated. The error incurred by the correlated rounding procedure can be bounded in terms of $|E_{\mathrm{adm}}|$, and the error from sampling can be bounded using concentration bounds.
2.2 Rounding Cluster LP
@@
In this section, we describe our algorithm for obtaining the improved approximation ratio for Correlation Clustering. We solve the cluster LP using Theorem 1 to get a fractional solution $z=(z_{S})_{S\subseteq V}$, which determines $x\in[0,1]^{V\choose 2}$ as in (2): $x_{uv}:=1-\sum_{S\supseteq\{u,v\}}z_{S}$ for every $uv\in{V\choose 2}$. We have $\mathrm{obj}(x)\leq(1+\varepsilon){\mathrm{opt}}$. The theorem will be proved in Section 4. With $z$, we then run two procedures: the cluster-based rounding and the pivot-based rounding with threshold $\tau=0.4$. We select the better result as the final clustering. The two procedures are defined in Algorithms 1 and 3 respectively. We use $N^{+}(u)$ and $N^{-}(u)$ to denote the sets of ++ and -neighbors of a vertex $u\in V$ respectively. They are stated as randomized algorithms, but they can be derandomized by the standard method of conditional expectation.
@@
Algorithm 2 General pivot-based rounding with hybrid independent/dependent updates. Parameters: $I_{c}^{+}\subseteq[0,1]$ and functions $f^{+},f^{-}:[0,1]\to[0,1]$. For a pivot $u$, ++edges $(u,v)$ with $x_{uv}\notin I_{c}^{+}$ are rounded independently (include $v$ with probability $1-f^{+}(x_{uv})$); ++edges with $x_{uv}\in I_{c}^{+}$ are rounded dependently via a sample $S\ni u$ drawn from $\{z_{S}\}$; -edges are always rounded independently (include $v$ with probability $1-f^{-}(x_{uv})$).
[…]
@@
3:  pick a pivot $u\in V^{\prime}$ uniformly at random;  $C\leftarrow\{u\}$
4:  sample a set $S\ni u$ according to probabilities $\{z_{S}\}$ $\triangleright$ $\sum_{S\ni u}z_{S}=1$
5:  for each $v\in V^{\prime}\cap N^{+}(u)$ do
6:   if $x_{uv}\notin I^{+}_{c}$ then $\triangleright$ independent rounding on out-of-range ++ edges
7:     independently add $v$ to $C$ with probability $1-f^{+}(x_{uv})$
8:   else if $v\in S$ then $\triangleright$ dependent rounding on in-range ++ edges
9:     add $v$ to $C$
10:  for each $v\in V^{\prime}\cap N^{-}(u)$ do
11:   independently add $v$ to $C$ with probability $1-f^{-}(x_{uv})$ $\triangleright$ independent rounding on - edges
12:  $\mathcal{C}\leftarrow\mathcal{C}\cup\{C\}$;  $V^{\prime}\leftarrow V^{\prime}\setminus C$
13:return $\mathcal{C}$
@@
Therefore, a ++edge uvuv will incur a cost of $\frac{2x_{uv}}{1+x_{uv}}$ in expectation in the cluster-based rounding procedure, and a -edge will incur a cost of $\frac{1-x_{uv}}{1+x_{uv}}$. The approximation ratios for a ++edge uvuv and a -edge uvuv are respectively $\frac{2}{1+x_{uv}}$ and $\frac{1}{1+x_{uv}}$. Notice that the latter quantity is at most $1$.
@@
We now proceed to the pivot-based rounding procedure in Algorithm 3.101010We remark that to recover the correlated rounding algorithm in [CLN22] and [CLLN23], we can use $C\leftarrow\emptyset$ in Step 4. Then we can obtain their approximation ratios without the complication of handling rounding errors. The errors are handled in [CLN22] by distinguishing between the short, median and long ++edges. In our algorithm, we also distinguish between short ++edges (those with $x_{uv}\leq\frac{1}{3}$) and long ++edges (those with $x_{uv}>\frac{1}{3}$); however, the purpose of this distinction is to get an improved approximation ratio, instead of to bound the rounding errors. Our high-level setup of the analysis is based on [ACN08] and [CMSY15] with some additional notions. We consider a general budget for every edge. We shall define two budget functions:
  • $\mathrm{budget}^{+}:[0,1]\to\mathbb{R}_{\geq 0}$ and $\mathrm{budget}^{-}:[0,1]\to\mathbb{R}_{\geq 0}$.
They determine the budget $b_{uv}$ for the edge $uv$: if $uv\in E^{+}$, then $b_{uv}:=\mathrm{budget}^{+}(x_{uv})$, and if $uv\in E^{-}$, then $b_{uv}:=\mathrm{budget}^{-}(x_{uv})$.
Algorithm 3 presents a generalized pivot-based rounding scheme parameterized by an interval $I_{c}^{+}\subseteq[0,1]$ and functions $f^{+},f^{-}$. Unlike the standard pivot rounding method, our scheme uses $I_{c}^{+}$ to designate the regime in which we apply dependent rounding (and applies independent rounding otherwise). By choosing different parameter triples $(I_{c}^{+},f^{+},f^{-})$, the algorithm instantiates a family of rounding procedures. We give two concrete instantiations that achieve different approximation guarantees: one more elementary and amenable to by-hand analysis, and one slightly more involved that yields a stronger approximation ratio.
We now focus on one iteration of the while loop in Algorithm 3. Suppose $u,v,w\in V^{\prime}$ at the beginning of the iteration, and let $C$ be the cluster constructed at the end. We use $u$ to denote the event that $u$ is chosen as the pivot. We say $vw$ incurs a cost in the iteration, if $vw\in E^{+}$ and $|C\cap\{v,w\}|=1$, or $vw\in E^{-}$ and $\{v,w\}\subseteq C$. Then we define
[…]
@@
$\frac{1}{|V^{\prime}|}\sum_{u\in V^{\prime}}\sum_{(v,w)\in{V^{\prime}\choose 2}}\mathrm{cost}_{u}(v,w)=\frac{1}{|V^{\prime}|}\sum_{(u,v,w)\in{V^{\prime}\choose 3}}\mathrm{cost}(u,v,w)+\frac{1}{|V^{\prime}|}\sum_{(u,v)\in{V^{\prime}\choose 2}}\mathrm{cost}(u,v).$
@@
To obtain an approximation ratio of $\alpha\in[1,2)$, we consider a variant of our algorithm, in which we run the cluster-based rounding procedure (Algorithm 1) with probability $\frac{\alpha}{2}$, and the pivot-based rounding procedure with threshold $\tau=0.4$ (Algorithm 3) with the remaining probability $1-\frac{\alpha}{2}$. Clearly, the actual algorithm that picks the better of the two clusterings generated can only be better. We set up the budget functions $\mathrm{budget}^{+}$ and $\mathrm{budget}^{-}$ such that every edge pays a cost of at most $\alpha$ times its LP cost in expectation. That is, the following properties are satisfied for every $x\in[0,1]$:
$\frac{\alpha}{2}\cdot\frac{2x}{1+x}+\left(1-\frac{\alpha}{2}\right)\mathrm{budget}^{+}(x)=\alpha x,\qquad\frac{\alpha}{2}\cdot\frac{1-x}{1+x}+\left(1-\frac{\alpha}{2}\right)\mathrm{budget}^{-}(x)=\alpha(1-x).$
@@
$\mathrm{budget}^{+}_{\alpha}(x):=\frac{\alpha}{1-\alpha/2}\cdot\frac{x^{2}}{1+x},\qquad\text{and}\qquad\mathrm{budget}^{-}_{\alpha}(x):=\frac{\alpha}{1-\alpha/2}\cdot\frac{(1+2x)(1-x)}{2(1+x)},\qquad\forall x\in[0,1].$ (5)
@@
If the budget functions $\mathrm{budget}^{+}_{\alpha}$ and $\mathrm{budget}^{-}_{\alpha}$ satisfy (4) for some $\alpha\in[1,2)$, then our algorithm has an approximation ratio of $\alpha$.